The satisfaction gained by consuming x units of good 1 and y units of good 2 is measured by the utility function




U = 2x2 + 5y3




Currently an individual consumes 20 units of good 1 and 8 units of good 2.




(a) Find the marginal utility of good 1 and hence estimate the increase in satisfaction gained from consuming one more unit of good 1.




(b) Find the marginal utility of good 2 and hence estimate the increase in satisfaction gained from consuming one more unit of good 2

1
Expert's answer
2022-02-01T06:55:39-0500

"a)Mu_1= 4x"

= "4\times 20= 80"

Satisfaction of consuming an extra unit of good 1

"= Mu_1\Delta1" ="80\times1= 80"

b) "Mu_ 2= 15Y^2"

"= 15(8^2)= 15\times 64= 960"

Satisfaction of consuming an extra unit of good

"=Mu_2\Delta2= 960\times 1= 960"